OpenStudy (anonymous):

Which of the points lie on the the line joining (2,0) to (4,3)? (a) (8,9) (b) (12,13) (c) (-4,-1) (d) (-6,-12) (e) (3 1/2, 2)

OpenStudy (anonymous):

what is the equation of the line passing through the two given points? which of those other five points are on that line?

OpenStudy (callisto):

equation of the line : y-0 3-0 ---- = ------ x-2 4-2 y = (3/2 )x - 3 Put the point into the equation to see if the equation exists.

jhonyy9 (jhonyy9):

- so i think e. is right - check it on diagramm

OpenStudy (callisto):

There should be 2 points satisfying the equation

OpenStudy (anonymous):

There are three points...

OpenStudy (anonymous):

Oh, I know how.

OpenStudy (anonymous):

can you find the slope of the line given those two points?

OpenStudy (callisto):

(y -y1) (y2-y1) ------ = ------- (x-x1) (x2-x1)

OpenStudy (anonymous):

or find the slope, then plug in one of the points to find the y intercept

OpenStudy (anonymous):

y= k(x-x1) + y1

OpenStudy (anonymous):

k : slope

OpenStudy (anonymous):

there are lots of ways to do everything... isnt it great?

OpenStudy (anonymous):

yea

OpenStudy (callisto):

Actually they are the same.. (y2-y1) k = ------- (x2-x1) Then you just multiply (x-x1) on both sides and +y1 on both sides

OpenStudy (anonymous):

right , vietnam?

OpenStudy (phi):

Which of the points lie on the the line joining (2,0) to (4,3)? I assume you can get the equation of the line through 2 points? y= (3/2) x -3 Now let's try the choices: a) (8,9) y= 12-3= 9 yes (b) (12,13) y= 18-3= 15 no (c) (-4,-1) y= -6-3= -9 no (d) (-6,-12) y= -9-3= -12 yes (e) (3 1/2, 2) y= 21/4 -12/4 = 9/4 no so choices (a) and (d)